OpenStudy (riley):

Does nobody know how to do this or at least explain it?!? I'm so confused uuughh. The latest demand equation for your Yoda vs. Alien T-shirts is given by: q = −40x + 800 where q is the number of shirts you can sell in one week if you charge $x per shirt. The Student Council charges you $400 per week for use of their facilities, and the T-shirts cost you $4 each. Find the weekly cost as a function of the unit price x.

OpenStudy (anonymous):

y=4x+400

OpenStudy (anonymous):

4$ a shirt and 400$ to rent the facilities, x=#shirts

OpenStudy (anonymous):

peace

OpenStudy (riley):

Thats incorrect >_<

OpenStudy (riley):

Is the same answer I've been getting, but its not right. Hence why I'm so confused, I don't get where I'm going wrong.

myininaya (myininaya):

Cost function=fixed cost + variable cost =400+(something) okay let's think we can only sell q shirts per week and each shirt cost us 4 dollars so i think the Cost function =400+4q =400+4(-40x+800) =400-160x+3200 =-160x+3600 but i'm not sure of the this terminology

OpenStudy (riley):

o_O

OpenStudy (riley):

Myininaya = my hero Thank you again :)

myininaya (myininaya):

np